1.An unfavorable variable overhead spending variance may be caused by:

A. the payment of lower prices for variable overhead items B. an increase in electricity costs per kilowatt hour charged by the utilities company C. price increases on direct materials D. the use of excessive quantities of variable overhead items

2.T Company expects to incur the following per unit costs for 1,000 units of production: Direct materials of 4 pounds per unit at $3 per pound AND direct labor of ¼ hour at $24 per hour AND variable overhead at 75 percent of direct labor costs AND fixed overhead of $3,000. What is the total amount of overhead to include in the overhead budget?

A. $13,500

B. $4,500

C. $3,000

D. $11,250

E. $7,500

3.V Company expected to sell 8,000 units at a price of $6 each. Instead, it sold 7,000 units at a price of $6.50 each. It hopes to sell even more next year, possibly 9,000 units. How much was the current period’s sales volume variance?

A. $6,500

B. $1,000

C. $6,000

D. $2,500

E. $500

4. Z Company is planning production of 200,000 units of Product X for June. One unit of X requires 2 pounds of raw material. The projected beginning inventory for June is 2,000 pounds. The projected ending inventory for June is 10,000 pounds. How many pounds of material should be purchased during June?

A. 392,000

B. 192,000

C. 208,000

D. 408,000

E. 416,000

5.H Company has the following expected pattern of collections on credit sales: 70 percent collected in the month of sale, 15 percent in the month after the month of sale, and 14 percent in the second month after the month of sale. The remaining 1 percent is never collected. April credit sales were $140,000. May credit sales were $160,000. Expected credit sales for June are $150,000. How much cash will H Company expect to collect in June?

A. $129,000

B. $152,520

C. $148,400

D. $127,400

E. $148,600
Answer #1

1. The answer is B.

2. Direct labor = 1/4 * $24 = $6

Total overhead = Variable overhead + Fixed overhead

= ($6 * 75%)* 1,000 + $3,000

= $7,500

The answer is E.

3. Sales volume variance = Flexible budget - Static budget

= (7,000 * $6) - (8,000 * $6)

= $6,000

The answer is C.

4. Raw material purchases = Units required for production + Ending inventory - Beginning inventory

= (200,000 * 2) + 10,000 - 2,000

= 408,000

The answer is D.

5. June cash collections = 70% of June sales + 15% of May sales + 14% of April sales

= ($150,000 * 70%) + ($160,000 * 15%) + ($140,000 * 14%)

= $148,600

The answer is E.